Please help me write and solve the equation for the word problem: Adayln has nine more than twice the number of skittles thsn Morgan has. If twice the number of skittles that Adalyn has is 2 more than 5 times the number of skittles that Morgan has, how many skittles does each girl have?
Answer by stanbon(75887)

Adayln has nine more than twice the number of skittles thsn Morgan has. If twice the number of skittles that Adalyn has is 2 more than 5 times the number of skittles that Morgan has, how many skittles does each girl have?
-----
A = 2M+9
2A = 5M+2
--------------------------------
Substitute for A and solve for "M"::
2(2M+9) = 5M+2
4M + 18 = 5M + 2
M = 16 (Morgan has 16 skittles)
-----
A = 2M+9
A = 2*16+9
A = 41 (Adayln has 41 skittles)
----------
Cheers,
Stan H.
----------